OpenStudy (anonymous):

Solve: 4(x - 1) - (x + 6) = 2 Answer 4 0 4 over 3 1

OpenStudy (anonymous):

x=4

OpenStudy (cwrw238):

4x-4-x-6 = 2 3x = 2+6+4 3x = 12 x=4